If 5 2 .5 4 .5 6 .5 8 .........5 2n =(0.04) -28 then n= OPTIONS: 5 6 7 8 9

If 52.54.56.58.........52n =(0.04)-28 then n=
OPTIONS:
  • 5
  • 6
  • 7
  • 8
  • 9

Grade:12th Pass

1 Answers

Sandeep Bhamoo
6 years ago
(0.04)^(-28) = (5^-2)^(-28) = 5^(56). Hence 56 = 2(1+2+3....n). 56=2{n(n+1)/2} =n(n+1). n=7. Please do approve my answer if correct